Find whether the two lines are perpendicular to each other.
y=5-3x 3x-y=8
Write an equation of the line containing the specified point and perpendicular to the indicated line. (0,-7), 4x+3-y

Answer by shree840(260)

Parallel lines have the same slope m
perpnedicular lines have slope in the ratio of m1*m2=-1
IT MEANS YOU FIND THE RECIPROCAL OF M AND CHANGE THE SIGN
Let us see
line1 y=-3x+5 m1=-3
line 2 3x-y=8 or y=3x-8 m2=3 m1*m2=3*-3=-9
line is not perpendicular